OpenStudy (anonymous):

The volumes of two similar solids are 729 m3 and 125 m3. The surface area of the larger solid is 324 m2. What is the surface area of the smaller solid? A) 56 m2 B) 100 m2 C) 500 m2 D) 200 m2 Thanks!

OpenStudy (anonymous):

i think it's A, just by logic.

OpenStudy (anonymous):

Nice! I wish I could reason like that!

OpenStudy (anonymous):

Hey Uzma, do you concure?

OpenStudy (anonymous):

volumes go with the cube, surface area with the square

OpenStudy (anonymous):

ratio is \[\frac{125}{729}\] which is the cube of \[\frac{5}{9}\]

OpenStudy (anonymous):

so the ratio of the surface areas will be \[\frac{25}{81}\]

OpenStudy (anonymous):

I'm following ya!

OpenStudy (anonymous):

and \[\frac{25}{81}\times 324=100\]

OpenStudy (anonymous):

that is your answer

OpenStudy (anonymous):

Thanks! its really helpful to see it written out like that!

OpenStudy (anonymous):

yw